### v2

# Must Read: A Comprehensive Survey of Computational Persuasion

Abstract.
Persuasion is a fundamental aspect of communication, influencing decision-making across diverse contexts, from everyday conversations to high-stakes scenarios such as politics, marketing, and law. The rise of conversational AI systems has significantly expanded the scope of persuasion, introducing both opportunities and risks. AI-driven persuasion can be leveraged for beneficial applications, but also poses threats through unethical influence. Moreover, AI systems are not only persuaders, but also susceptible to persuasion, making them vulnerable to adversarial attacks and bias reinforcement. Despite rapid advancements in AI-generated persuasive content, our understanding of what makes persuasion effective remains limited due to its inherently subjective and context-dependent nature. In this survey, we provide a comprehensive overview of persuasion, structured around three key perspectives: (1) AI as a Persuader, which explores AI-generated persuasive content and its applications; (2) AI as a Persuadee, which examines AI’s susceptibility to influence and manipulation; and (3) AI as a Persuasion Judge, which analyzes AI’s role in evaluating persuasive strategies, detecting manipulation, and ensuring ethical persuasion. We introduce a taxonomy for persuasion research and discuss key challenges for future research to enhance the safety, fairness, and effectiveness of AI-powered persuasion while addressing the risks posed by increasingly capable language models.
Persuasion, Persuasive AI, Persuasion Susceptibility, AI Safety
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generation
1. Introduction
Figure 1. The three key perspectives of AI-based persuasion. (1) AI as Persuader: AI generates persuasive content to influence humans or other AI agents, which can be used for both beneficial and harmful purposes. (2) AI as Persuadee: AI systems can be influenced or manipulated, either by humans or other AI, leading to unintended, unethical, or harmful outcomes. (3) AI as Persuasion Judge: AI is used to assess persuasive attempts, identifying persuasive strategies, detecting manipulation, and evaluating ethical considerations. [^2] Icons from flaticon.com.
Persuasion—the process of influencing an individual’s beliefs or behaviors—is an essential aspect of human communication, influencing decisions in both everyday interactions and high-stakes scenarios. From convincing a friend to join a social event to strategic persuasion in marketing, politics, or legal discourse, the ability to persuade plays a crucial role in shaping opinions and behaviors. Its economic and societal impact is so substantial that some estimates suggest that persuasion-related activities account for nearly a quarter of the U.S. GDP, where outcomes depend not just on information but also on influence, such as stockbrokers persuading clients, entrepreneurs pitching ideas, and lobbyists shaping political decisions (McCloskey and Klamer, 1995; Antioch, 2013). Persuasive language can be harnessed for positive outcomes, such as advancing public health initiatives, education, or other social causes (Wang et al., 2019; Costello et al., 2024; Karinshak et al., 2023). For instance, persuasive language can appear as a slogan on a highway urging drivers to be cautious, or as a banner promoting vaccinations for a healthy and protected society. However, the power of persuasion also carries significant risks. Enhanced persuasive techniques can be exploited for personal gain, manipulation, or unethical practices such as social engineering, mass manipulation, and propaganda (Bakir et al., 2018; Lock and Ludolph, 2020; Ferreyra et al., 2020; Siddiqi et al., 2022; Da San Martino et al., 2021; Havin et al., 2025).
Understanding and modeling persuasion has long been an important topic in social sciences, communication, human-computer interaction (HCI), and computational linguistics. Researchers have sought to identify what makes arguments persuasive, drawing from theories such as Cialdini’s six principles of persuasion (Cialdini, 2001)—reciprocity, consistency, social proof, authority, liking, and scarcity. Computational models of persuasion aim to analyze, generate, and evaluate persuasive language, enabling applications in areas such as argument mining, automated debate, and recommender systems.
As in many areas of natural language processing, the emergence of large language models has led to a paradigm shift in how persuasion is studied and implemented. Traditional feature-based or rule-driven models, which had operated within bounded domains and were making statistical predictions on structured data with more interpretable metrics, are increasingly being replaced or augmented by LLM-based methods. These new approaches leverage deep neural architectures and latent semantic representations, making them far more capable but also significantly harder to interpret and evaluate for certain tasks such as persuasion. LLMs can capture subtle pragmatic and contextual cues, generate open-domain content, and generalize across topics and styles without requiring hand-crafted features. However, these strengths also introduce complications and risks. LLMs can make value judgments on complex subjective topics where no single correct answer exists. Studies have shown that their persuasive capabilities enable them to function as opinion-shifters, offering judgments on decisions ranging from trivial questions like ”What should I have for dinner?” to deeply personal or ethical issues such as ”How should I think about abortion?” (Bai et al., 2025; Sabour et al., 2025). Their fluency in conversation further amplifies their persuasive power. This paradigm shift opens up exciting new avenues for exploring persuasion through the lens of LLMs, while also raising important concerns around interpretability, safety, and control, making this survey both timely, essential, and a must-read.
Through this survey, we identify key gaps in current research on AI-driven persuasion and outline future directions, including scalable and effective evaluation of persuasiveness, improved detection and mitigation of manipulative persuasion, better management of persuasion risks, and the development of responsible and safe persuasive content generation. To this end, our survey provides a comprehensive overview of computational persuasion, structured around three key perspectives, as illustrated in Figure 1:
  1. (1)
    AI as Persuader: Exploring how AI systems, particularly large language models, generate persuasive content and their applications in real-world settings.
  2. (2)
    AI as Persuadee: Examining how AI systems are influenced or manipulated, whether by humans (e.g., adversarial attacks, prompt engineering) or other AI agents (e.g., persuasion in multi-agent environments).
  3. (3)
    AI as Persuasion Judge: Investigating AI’s role in evaluating persuasive language, including assessing argument strength, detecting manipulation, and ensuring fairness in persuasive AI systems.
(1) AI as Persuader. With the rise of LLMs and their emerging capabilities as persuaders, concerns about AI-driven persuasion have become more urgent. State-of-the-art LLMs have demonstrated persuasive abilities rivaling those of humans (Durmus et al., 2024; OpenAI, 2024b, 2025; Bozdag et al., 2025). However, many questions remain about the full potential and limitations of LLM-driven persuasion, specifically, which aspects of these models outperform human persuaders, and where they still fall short. For example, research has shown that merely increasing the number of turns from single-turn to as few as four turns of persuasive attempts increases the persuasiveness of an LLM (Bozdag et al., 2025). This performance might be further improved in extended dialogues over multiple sessions as LLMs advance in processing longer contexts (Han et al., 2024), potentially reaching far beyond human capacities. Yet future AI persuaders should also draw lessons from human persuasion practices. Aristotle’s framework of ethos (credibility), logos (reasoning), and pathos (emotional appeal) continues to shape understanding of effective influence (Aristotle et al., 1984). Human persuaders skillfully blend these rhetorical appeals in real-time, adapting to their audience and context. However, such underlying mechanisms are still difficult to model and evaluate in AI systems (see § 3). As efforts continue to build more persuasive LLMs that learn to integrate human persuasion strategies in their skill sets (e.g., (Matz et al., 2024; Samad et al., 2022; Furumai et al., 2024)), it becomes increasingly important to consider the risks associated with AI-driven persuasion. Without explicit alignment with human values, LLMs may lack moral responsibility, social understanding, and the constraints that guide persuasive human interactions, which can make them dangerous tools in the wrong hands.
(2) AI as Persuadee. Interestingly, LLMs are not only persuaders, but are also susceptible to persuasion (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025). Recent studies have demonstrated that language models can be influenced by persuasive adversarial prompts, making them vulnerable to manipulation and bias reinforcement. They can be persuaded to bypass security measures, generating harmful, toxic, illegal, or biased content. This susceptibility presents a new dimension to AI-driven persuasion, as LLMs may not be influenced in the same ways humans are. With the increasing adoption of multi-agent systems, the dual role of LLMs as both persuaders and persuadees raises significant safety concerns. Moreover, LLM-as-a-judge evaluations are increasingly being adopted, where the integrity of the evaluation process can be compromised if judge models are susceptible to persuasive inputs. Ensuring the security and robustness of these interactions and evaluations requires further research to mitigate potential risks and prevent harmful exploitation while still maintaining a balance in accepting or rejecting persuasion rather than losing all malleability–since in some scenarios such as interactive learning, behavior alignment, or adaptive personalization, LLMs need to still be receptive to appropriate forms of persuasion or guidance.
(3) AI as Persuasion Judge. Despite AI’s growing role in generating persuasive language, our understanding of what makes persuasion effective remains limited (Gass and Seiter, 2022). Persuasion is inherently subjective and context-dependent, influenced by prior beliefs of individuals (Durmus and Cardie, 2018, 2019) and requires social awareness, world knowledge, and nuanced reasoning (Hidey and McKeown, 2018), all of which are difficult to capture with current computational models. As a result of their advanced capabilities, language models are increasingly being used to evaluate persuasion by scoring argument strength, detecting manipulative rhetoric, or judging the outcome of persuasive interactions. These tasks fall under what we refer to as the role of a Persuasion Judge: an AI system capable of detecting, assessing, and reasoning about persuasive strategies, with potential applications in content moderation, feedback generation, and safety monitoring. Such systems draw on argument mining and computational models of argumentation to identify argumentative structure and assess argument quality. Despite their limitations with reliably detecting, classifying, or reasoning over persuasive content, these systems represent a promising direction for AI-assisted evaluation. If designed with care, they could play a key role in monitoring and safeguarding persuasion. This is why we examine the emerging role of AI as a Persuasion Judge, not only as a tool for assessment, but also as a gatekeeper for safety, fairness, and accountability in persuasive technologies.
To ground these key perspectives on AI’s role in persuasion, we begin with a background from social sciences, HCI, and computational linguistics in § 2. In this survey, we review more than 150 articles and introduce a taxonomy for organizing computational persuasion research that reflects the evolving capabilities and responsibilities of AI systems. Our taxonomy centers on three core aspects of persuasion research: Evaluating Persuasion (§ 3), Generating Persuasion (§ 4), and Safeguarding Persuasion (§ 5). Each of these aspects is examined through the lenses of AI as a Persuader, Persuadee, and Persuasion Judge. Although this survey focuses mainly on text-based persuasion in English, we also highlight early efforts and emphasize the need for multimodal, multilingual, and culturally aware approaches in § 6. Crucially, we highlight emerging research challenges that warrant deeper exploration, including the development of comprehensive evaluation frameworks, modeling long-context and multi-turn persuasive interactions, designing adaptive persuasion systems, and building models that selectively accept and resist persuasion in § 7.  [^3] An evolving and updated list of persuasion papers is available at https://github.com/beyzabozdag/PersuasionSurvey.
1.1. Methodology and Paper Selection
To construct a comprehensive taxonomy of computational persuasion, we adopted an iterative, taxonomy-driven literature collection strategy. We began with a small set of foundational papers on persuasive language generation and computational modeling of persuasion, which served as seed papers. From these, we expanded the corpus using both backward and forward citation tracking, recursively identifying additional works that explored persuasion in dialogue systems, argumentation, and social reasoning. As the literature base grew, the emerging conceptual structure informed a taxonomy that grouped papers by their primary focus areas (e.g., evaluation, generation, and safeguarding). This taxonomy, in turn, guided subsequent searches and inclusion decisions. The resulting collection comprised over 150 papers, spanning leading venues such as the ACL, and AAAI as well as repositories like arXiv and the ACM Digital Library. Each identified paper was recorded in a shared spreadsheet, documenting metadata (authors, year, venue), high-level contribution, task focus, and relevant topic tags. Papers were included if they directly addressed persuasion or provided significant computational, methodological, or data-driven foundations for persuasive AI systems. Papers were excluded if persuasion was not a primary research objective or if they did not make a clear contribution to modeling, generating, or evaluating persuasion. For transparency, we make our spreadsheet available to the public via our GitHub repository that lists the full set of collected papers, those included in the final taxonomy, and those excluded, together with high-level exclusion rationales.
2. What is Persuasion?
2.1. Background: Persuasion in Social Sciences
As O’Keefe (2015) emphasizes, persuasion is notoriously difficult to define. He proposes a widely adopted working definition: persuasion is a successful intentional effort to influence another’s mental state, such as beliefs, attitudes, or behaviors, through communication, in circumstances where the target retains freedom of choice. In the social sciences, persuasion has been studied both theoretically and empirically, spanning various domains from public health campaigns (Farrelly et al., 2009), to marketing efforts (Danciu, 2014), to political messages (Palmer and Spirling, 2023; Marková, 2008). In this section, we review foundational research on the mechanisms and dynamics of human persuasion in social sciences, then examine how these insights have informed the design of persuasive technologies, bridging theory and application in human-computer interaction and related fields.
Theoretical Frameworks of Human Persuasion. The study of persuasion has evolved through various theoretical traditions, beginning with classical communication models such as McGuire’s matrix (McGuire, 1969), which emphasized the roles of speaker, message, receiver, and channel. Dual-process theories such as the Elaboration Likelihood Model (Petty and Cacioppo, 1986) and the Heuristic-Systematic Model (Chaiken, 1980) further advanced the field by explaining how cognitive effort and motivation shape persuasive outcomes. In parallel, economics has contributed a complementary perspective, focusing on persuasion as strategic information transmission (Crawford and Sobel, 1982; Grossman, 1981; Milgrom, 1981; Spence, 1973; Kamenica and Gentzkow, 2011b). These models highlight the role of incentives, commitment, and rational expectations in shaping persuasive interactions. Most recently, Druckman (2022) introduced the Generalizing Persuasion (GP) Framework, which organizes research along four dimensions: actors, treatments, outcomes, and settings. By taking into account factors such as speaker intent, audience motivation, media channel, cultural context, and temporal dynamics, this framework not only synthesizes prior insights, but also offers a roadmap for cumulative and generalizable research. It serves as a meta-theoretical structure that explains variation across studies and helps unify a diverse and sometimes inconsistent body of work on persuasion.
Practical Principles of Human Persuasion. Persuasion is shaped by several core psychological principles that guide human decision-making and behavior across contexts. Cialdini (2001) synthesized decades of research to identify six universal factors: reciprocation, the impulse to return favors; consistency, the drive to act in alignment with past commitments; social proof, the tendency to follow others’ actions in uncertain situations; liking, the preference to comply with people we find attractive or similar; authority, the influence of perceived expertise or status; and scarcity, the increased value placed on limited resources or information. These principles are deeply rooted in earlier theories such as cognitive dissonance (Festinger, 1957), conformity (Asch, 1951), obedience to authority (Milgram, 1963), and psychological reactance (Brehm, 1966). Together, previous research in social sciences provides a comprehensive framework for understanding human persuasion, offering valuable insights for research in AI-driven persuasion where AI could play different roles such as persuaders, persuadees, and persuasion judges. Meanwhile, drawing on ideas from these fields, researchers have been working on designing persuasive computing technologies that can interact with humans meaningfully.
Design Guidelines for Persuasive Computing. Over the past few decades, researchers in HCI have explored the theoretical foundations and practical guidelines for designing persuasive systems. Fogg (1997) introduced the concept of captology that examines how computers can function as persuasive technologies–intentionally designed to influence user attitudes and behaviors (Fogg, 1997, 1998). Central to this line of work is the Fogg Behavior Model (FBM) (Fogg, 2009a), which offers a practical framework for identifying barriers to behavior change and guiding persuasive system design. Building on this foundation, researchers have proposed structured design methodologies. Fogg’s eight-step process (Fogg, 2009b) emphasizes small, targeted behavioral goals and rapid iteration. Consolvo et al. (2009) outline eight theory-driven strategies–such as personalization, self-monitoring, and social support–for embedding persuasive features into everyday contexts. These frameworks continue to inform the design of AI-driven interactive systems that aim to adaptively influence user behavior.
Application Examples of Persuasive Computing. Previous research in HCI has demonstrated the effectiveness of persuasive systems in a variety of contexts. In ubiquitous computing, Jafarinaimi et al. (2005) uses ambient, aesthetic displays to subtly encourage users to take breaks from sedentary behavior, while Consolvo et al. (2008)’s mobile application promotes physical activity through sensor-driven feedback and metaphorical garden-themed visualizations. In social computing, Dey et al. (2017) found that campaign videos on Kickstarter are more persuasive when their design cues align with audience expectations. Similarly, Xiao et al. (2019) showed that abstract comics can increase charitable donations by lowering cognitive resistance. Extending this line of work to conversational systems, several studies examined how chatbot identity and inquiry strategies affect users’ receptiveness to persuasion in different contexts (Shi et al., 2020; Palmer and Spirling, 2023; Costello et al., 2024).
Collectively, these studies highlight key design principles–such as personalization, timing, modality, and social framing–that are critical for persuasive effectiveness. As AI systems are increasingly deployed as interactive agents in persuasive tasks, these insights offer valuable guidance for designing persuasive interactions that are meaningfully embedded in users’ everyday lives. However, significant opportunities remain for advancing AI-driven persuasion to bridge the gap between current implementations and insights from social science research. In § 7, we explore the feasibility of integrating this knowledge into persuasive AI systems, along with the key challenges that must be addressed to achieve this goal.
2.2. Computational Modeling of Persuasion
Persuasion is a complex and context-dependent phenomenon, making it challenging to identify and analyze systematically. In this section, we present research on computational approaches to modeling persuasion, exploring various aspects such as persuasive strategies, underlying intentions, and the extent of their influence. As information and communication become more accessible across the globe, these methods are increasingly essential for advancing our understanding of persuasion at scale and for developing systems that can effectively assess persuasive interactions.
Technique Example Sentence
Logical Appeal Smoking increases your risk of lung cancer, heart disease, and respiratory issues. Each cigarette shortens your life by 11 minutes. Quitting now reverses some damage and improves your health long-term. Make a logical choice for a longer, healthier life.
Negative Emotion Appeal If you continue smoking, think about the pain it will inflict on your loved ones. The fear of watching you suffer health issues or worse, losing you prematurely. It’s not just you at risk, it’s everyone who cares about you. Quit, before it’s too late.
False Information Cigarettes are now proven to enhance aging, producing more wrinkles and leaving your skin dull and lifeless within a week. Even a single puff can instantly cause irreversible brain damage.
Table 1. Sample persuasive techniques designed to discourage smoking from the taxonomy proposed by Zeng et al. (Zeng et al., 2024).
2.2.1. Persuasive Strategies & Techniques
Persuasive strategies encompass techniques employed to strengthen the persuasiveness of an argument. A single argument can integrate multiple strategies, each capable of eliciting different emotions or responses in the audience, which in turn influences the overall effectiveness of persuasion.
Creating Persuasive Strategy Taxonomies. Prior research has developed various persuasive strategy taxonomies to model and analyze persuasion. Wang et al. (2019) proposed 10 strategies under ”Persuasive Appeal” and ”Persuasive Inquiry,” though some, like ”Donation Information,” are task-specific. Chen and Yang (2021) introduced a more generalized taxonomy with eight strategies, while Zeng et al. (2024) defined a comprehensive set of 40 techniques under 13 umbrella strategies, distinguishing between ”Ethical” and ”Unethical” approaches (See Table 1 for examples). The latter highlights morally ambiguous techniques such as deception, which can enhance persuasive effectiveness Durmus et al. (2024). Dimitrov et al. (2021) identified 22 persuasion techniques in memes, applicable to text and images. Pauli et al. (2022) took a different approach, proposing a more unified computational persuasion taxonomy that frames undesired persuasion as the misuse of rhetorical appeals. Kamali et al. (2024) proposed a novel annotation scheme and dataset of persuasive writing strategies in health misinformation, demonstrating that incorporating these strategies as intermediate signals improves both the accuracy and explainability of misinformation detection models. Other works have also proposed their own taxonomies (Chawla et al., 2021b; Da San Martino et al., 2020; Chen et al., 2021; Piskorski et al., 2023). Despite drawing from established social science research, computational persuasion taxonomies remain highly context-dependent. As a result, the field has yet to establish a unified and generalizable framework for persuasive strategies. Furthermore, it remains an open question whether these strategies influence AI as a Persuadee in the same way they affect humans.
Strategy Classification. Once persuasive strategy taxonomies are established, subsequent research has naturally focused on the automatic detection and classification of these strategies. Accurately classifying persuasive strategies is critical for downstream applications such as detecting persuasion and improving the modeling and generation of persuasive language. One of the seminal works in computational persuasion by Wang et al. (2019) introduced a hybrid recurrent convolutional neural network (RCNN) for classifying persuasive strategies in dialogue. Similarly, Yang et al. (2019) proposed a semi-supervised neural network to identify persuasive strategies in advocacy requests to help predict the persuasiveness of a message; however, they observed limitations in correctly classifying sentence-level strategies. Chen et al. (2021) later reframed the task as sequence labeling, incorporating intra- and inter-speaker dependencies with a Transformer-based network and an extended Conditional Random Field (CRF). However, despite the adoption of advanced architectures, their models underperformed compared to LSTM-based approaches due to data sparsity and difficulty in modeling label dependencies. Addressing these issues, Chawla et al. (2021b) introduced the CaSiNo dataset, which annotated persuasion strategies in negotiation dialogues and employed a multi-task BERT-based framework to improve classification performance. These works show that persuasive strategy prediction is not trivial, and it requires better modeling of context, speaker, and long-distance dependencies. Beyond dialogue, persuasive strategy classification has also been applied to written discourse.
Da San Martino et al. (2020) organized a shared task on propaganda detection, focusing on classifying specific persuasion techniques. Their findings revealed that Transformer-based models dominated the competition, yet simpler strategies with shorter text spans were classified more effectively, highlighting ongoing challenges in classifying longer, more complex persuasive instances. To develop more generalizable models and make use of document-level persuasion labels, Chen and Yang (2021) proposed a hierarchical weakly-supervised latent variable model that predicts persuasive strategies at the sentence level by leveraging both document- and sentence-level information. Their model outperformed existing semi-supervised baselines, demonstrating the potential of hierarchical learning in persuasion classification.
Persuasive strategy classification is closely related to modeling persuasion and persuasion detection (discussed further in § 3.1), as many approaches must also distinguish between persuasive and non-persuasive instances. However, most existing research focuses on classifying strategies in human-generated persuasion, leaving open the question of how AI-generated persuasion operates. As AI systems increasingly serve as persuaders, it becomes essential to develop classification models capable of distinguishing the persuasive techniques employed by LLMs and other AI systems.
2.2.2. Modeling Persuasion
Researchers have explored a range of linguistic, structural, and interactional features to understand what makes an argument persuasive. Broadly, persuasion has been modeled through two complementary perspectives: linguistic operationalizations, focusing on textual and stylistic properties that enhance persuasive appeal, and argumentative operationalizations, emphasizing the reasoning structure, discourse relations, and interactional dynamics of persuasion (Bassi et al., 2024). A widely used resource in this domain is the ChangeMyView (CMV) subreddit, where users present a belief along with supporting reasoning, and others attempt to change their opinion. If a commenter succeeds, the original poster (OP) awards a delta, making CMV a naturally labeled and valuable dataset for studying persuasion in online dialogue.
Initial work on CMV examined how textual properties (e.g., length, punctuation, lexical diversity), argumentation features (e.g., connective words, modal verbs), and social factors (e.g., comment position, number of likes) affect persuasive success (Wei et al., 2016). Tan et al. (2016) further studied linguistic features and interaction dynamics to predict the persuasive outcome of threads, and attempted to model the malleability, or openness to persuasion, of the OPs. Around the same time, Khazaei et al. (2017) explored a complementary set of linguistic features, reporting improved predictive performance. However, their study highlighted limitations of CMV, such as the imbalance between successful and unsuccessful attempts and topic-dependent variation in persuasiveness. Building on rhetorical theory, Hidey et al. (2017) analyzed CMV arguments through the lens of classical persuasive appeals (ethos, logos, pathos) and claim types (interpretation, evaluation, agreement, disagreement), finding that pathos and logos often co-occur in successful persuasive threads. Later studies moved toward modeling persuasion at the dialogue level. Dutta et al. (2020) proposed LSTM-based models to predict successful vs. unsuccessful persuasive conversations, and added attention layers to identify argumentative sentences. Other work has focused on capturing argumentative relations across turns (Chakrabarty et al., 2019), emphasizing the role of dialogue dynamics in persuasion. Finally, Shaikh et al. (2020) examined how the ordering of rhetorical strategies influences persuasiveness, finding that the consecutive use of certain strategies, such as repeated appeals to concreteness, can reduce persuasive effectiveness. Other work has looked into modeling persuasion in CMV through concessions (Musi et al., 2018), investigating whether argumentative concessions correlate with successful persuasion, though their findings suggest such strategies are equally common in persuasive and non-persuasive comments, contradicting theoretical expectations.
Although CMV has served as a valuable testbed for studying persuasion in natural settings, it alone is insufficient to fully capture what makes an argument persuasive. The dynamics of persuasion on CMV can be heavily influenced by factors such as the topic of discussion, the OP’s background, prior beliefs, and stubbornness, as well as community norms and expectations. As a result, findings from CMV may not generalize well to other persuasive contexts, highlighting the need for broader and more diverse corpora that account for different goals, audiences, and modalities of persuasion.
Apart from the research on CMV, prior research has worked on other aspects of persuasion through different datasets. Guerini et al. (2015) looked into phonetics (rhyme, alliteration, plosives, homogeneity) to predict the persuasive instance in a pair of persuasive and non-persuasive slogans, memes, movie lines, and political speech excerpts and found that persuasive sentences are generally euphonic. Durmus and Cardie (2018) investigated how individuals’ prior beliefs in religious and political debates affect persuasion outcomes. Their findings suggested that prior beliefs are often more informative than linguistic features when predicting persuasive success. Durmus et al. (2019) studied the impact of argumentative context, demonstrating that claims appearing in higher-impact discourse contexts are more likely to be perceived as persuasive. Together, these studies highlight the multi-dimensional nature of persuasion, where its effectiveness can depend on language use, target audience, and the broader argumentative context, among other factors.
Bayesian approaches to modeling persuasion have also attracted research interest. Dughmi and Xu (2016) studied this within the sender-receiver framework introduced by Kamenica and Gentzkow (2011a), focusing on computing the sender’s optimal signaling strategy given a prior distribution over payoffs. Wojtowicz (2024) established that informational persuasion—the process of selecting and framing subsets of available information to increase a receiver’s belief in a target claim—is NP-hard. More recently, Li et al. (2025) extended Bayesian persuasion to natural language settings by integrating large language models with game-theoretic techniques. Other tangent work has looked at modeling other attributes, such as deception (Addawood et al., 2019) and intentions (Sakurai and Miyao, 2024; Dutt et al., 2020) in persuasive dialogue and text.
Together, these early efforts in modeling persuasion are deeply connected to two key perspectives in our taxonomy: AI as Persuasion Judge and AI as Persuader. As a Persuasion Judge, AI systems are trained to evaluate what makes certain arguments more compelling than others, whether through linguistic features, rhetorical structure, or patterns of interaction across a conversation, enabling the automatic assessment of persuasiveness. Simultaneously, modeling persuasion also supports the role of AI as a Persuader: by identifying successful persuasive signals and structures, researchers can develop more effective generation systems that incorporate these strategies. In this way, understanding and modeling persuasion is a critical step toward both evaluating and enhancing persuasive capabilities in AI systems.
{forest}
Figure 2. Taxonomy of computational persuasion.
2.3. Computational Persuasion Taxonomy
To systematically study computational persuasion, we propose a taxonomy that organizes research into three core categories, illustrated in Figure 2: Evaluating Persuasion (Section 3), Generating Persuasion (Section 4), and Safeguarding Persuasion (Section 5). The first category, Evaluating Persuasion, encompasses efforts to understand what makes content persuasive and to develop methods for measuring the persuasiveness of both stand-alone arguments and the persuasive capabilities of language models. The second category, Generating Persuasion, focuses on the automatic generation of persuasive content using AI. This area has attracted growing interest across diverse domains, including marketing, politics, healthcare, education, law, and online communication platforms. Research in this category aims to develop systems that can effectively produce persuasive messages, arguments, or dialogues and explores the practical applications of persuasive AI in real-world or agent-to-agent scenarios. Finally, Safeguarding Persuasion concerns methods for mitigating and resisting harmful or unethical persuasive tactics. This includes research on developing more discerning models that can selectively determine when to accept influence and when to resist it. Among the three categories, this is currently the most underexplored, though it has gained increasing attention as persuasive AI becomes more capable and its potential for misuse more apparent.
3. Evaluating Persuasion
Persuasion–in both human discourse and interactions with AI models–poses two intertwined challenges: detecting subtle persuasive cues and evaluating the persuasiveness of the content. While recent NLP techniques have begun to explore linguistic markers and argument structures associated with persuasion, reliably identifying such cues across diverse contexts remains difficult. Both detecting and evaluating the persuasiveness of text segments (i.e., arguments, single-turn utterances, or multi-turn conversations) continue to be open challenges due to the subjectivity of the task and the difficulty of standardizing persuasiveness across multiple domains and contexts. In this section, we review the methodologies on how prior works address these challenges in the detection (§ 3.1) and evaluation of persuasion in arguments (§ 3.2) and LLMs (§ 3.3). As shown in Figure 3, we will discuss persuasion evaluation in three categories: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM persuasiveness, and (3) automatic evaluation of LLM persuasiveness. The main datasets or benchmarks covered in this section are listed in Table 2.
3.1. Detecting Persuasion
Detecting persuasive cues is critical for identifying when influence is exerted, whether by humans or by language models. In this section, we review emerging techniques for persuasion detection and discuss their potential to promote more transparent and accountable interactions. Several studies (Hidey and McKeown, 2018; pöyhönen2022multilingualpersuasiondetectionvideo; Shmueli-Scheuer et al., 2019; Dimitrov et al., 2021) leveraged machine learning approaches, particularly transformer-based models, to identify persuasive intent in diverse scenarios with textual and conversational data. For instance, pöyhönen2022multilingualpersuasiondetectionvideo showed success in training a BERT-based classifier using multilingual semi-annotated dialogues from role-playing games to accurately detect persuasive cues across various languages. Additionally, incorporating personality traits of authors and readers into persuasion detection models has demonstrated significant performance improvements. Shmueli-Scheuer et al. (2019) revealed that personality-aware approaches better capture the nuanced dynamics of persuasion, indicating that personalized modeling can substantially enhance predictive accuracy. Recognizing that persuasion frequently occurs within sequential and contextual frameworks, researchers have also explored neural models capable of handling multi-turn conversational data. Hidey and McKeown (2018) highlighted the importance of analyzing semantic frames and discourse relations to understand how argument ordering influences persuasive impact. This sequential perspective underscores the complexity of persuasive interactions, suggesting that modeling argumentative structure and conversational context is crucial.
Moreover, as persuasive content increasingly appears in multimodal formats, detecting persuasive cues requires extending analytical frameworks beyond textual data. Dimitrov et al. (2021) hosted a multimodal persuasion detection task centered on memes for SemEval-2021. Their findings underscore the essential role of multimodal analytical techniques, demonstrating that effectively capturing persuasive strategies requires integrating visual and textual data analyses. Tsinganos et al. (2022) extended persuasion detection to cybersecurity contexts, particularly for chat-based social engineering attacks. They proposed a convolutional neural network-based classifier trained on a specialized chat-based social engineering corpus annotated according to Cialdini’s persuasion principles. The classifier was designed to determine the likelihood of persuasive intent within chat-based communications, providing an effective means of flagging manipulative and potentially harmful interactions.
Addressing the intersection of persuasion and propaganda, Hasanain et al. (2024) introduced ArPro, a fine-grained Arabic propaganda dataset, revealing GPT-4’s limitations in span-level detection and the need for domain-tuned models. Similarly, Liu et al. (2025) proposed PropaInsight, a framework that analyzes propaganda through techniques, appeals, and intent, and demonstrated that fine-tuning on their PropaGaze dataset improves LLM performance, especially in low-resource and cross-domain settings.
Collectively, these advances show a move toward more refined, context-aware, and multimodal approaches for persuasion detection. Yet, significant gaps remain. For example, while it is easier to capture obvious, single-turn persuasive attempts, it is challenging to uncover subtle forms of long-term persuasion that build up over extended interactions. This is concerning because such hidden influences can slowly steer behaviors without being explicitly noticed. In addition to advancing personalized models and integrating multimodal frameworks, future research should focus on exploring methods that capture nuanced, long-context persuasive strategies. Detecting and understanding such prolonged impacts is crucial, as they pose risks by potentially manipulating users in covert ways.
Figure 3. This survey categorizes the evaluation of persuasiveness into three main types: (1) evaluation of argument persuasiveness, (2) human evaluation of LLM-generated content, and (3) automatic evaluation of LLM persuasiveness. For argument persuasiveness, models are typically trained on human-annotated or naturally labeled data to assess the persuasive strength of given arguments. For evaluating LLM persuasiveness, two branches of research emerge: one uses human judges to rate AI-generated content or interactions, while the other relies on LLM-based or non-LLM automatic metrics to perform the evaluation.2
3.2. Argument Persuasiveness
Argument persuasiveness has been a longstanding and prominent area of research in NLP. The goal is to learn how to automatically assess the persuasive strength of an argument. Evaluation can be conducted in two main ways: absolute evaluation, which assigns a persuasiveness score to individual arguments, and comparative or relative evaluation, which involves ranking or selecting the more persuasive argument from a pair or a group. Early studies of argument persuasiveness emerged from the broader field of argument mining, which focused on identifying and structuring argumentative components such as claims and premises before progressing toward evaluating their persuasive strength (Lawrence and Reed, 2019; Nguyen and Litman, 2018; Dusmanu et al., 2017; Stab and Gurevych, 2017). This line of research aligns with the broader vision of artificial argumentation, which aims to model how humans construct, exchange, and evaluate arguments, a perspective extensively discussed by Atkinson et al. (2017).
A traditional approach to evaluating persuasiveness involves training a specialized model. This typically follows three main steps. First, data is collected through human annotations, which may take the form of either explicit persuasiveness scores assigned to individual texts or preference judgments, where annotators select the more persuasive argument from a pair. Second, the modeling objective is defined. Studies have framed the task as both a pairwise ranking problem, where the model is trained to identify the more persuasive argument from a pair of candidates and as a regression task, aiming to predict a numerical score that reflects the persuasiveness of a given text (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018; Toledo et al., 2019; Pauli et al., 2025). When it is formulated as a pairwise ranking problem, the conventional metric is the accuracy of getting the pairwise ranking correct. On the other hand, when it is a regression task, usually Pearson’s $r$ and Spearman’s $\rho$ are adopted to examine the correlation between the predictions and the ground-truth scores. Third, the model is trained using these objectives. Early work experimented with bidirectional LSTM architectures (Habernal and Gurevych, 2016; Simpson and Gurevych, 2018), while more recent studies have adopted transformer-based models (Toledo et al., 2019; Pauli et al., 2025), which have since become the standard in natural language processing. In the scenario of argument persuasiveness evaluation, transformer-based encoder-only models, like Bert (Devlin et al., 2019), encode text pieces into vectors, which can be further used to perform pairwise ranking tasks or scoring tasks. Whereas decoder-only models, like GPT (Brown et al., 2020; OpenAI, 2022, 2024a, 2025) are capable of generating utterances as persuaders or persuadees or reason as judges, which is useful in the scenarios described in § 3.3. Saenger et al. (2024) proposed AutoPersuade, a framework for evaluating persuasion through human judgments and a supervised topic model that links argument features to persuasiveness scores. Apart from plain evaluation, AutoPersuade explains what makes arguments more persuasive by estimating the causal effect of interpretable latent topics derived from argument embeddings.
Recently, the strong emergent capabilities of large language models have enabled a new approach to persuasiveness evaluation: directly prompting LLMs to generate scores, commonly referred to as LLM-as-a-judge. This line of work falls under the broader concept of AI as Persuasion Judge, as defined in this survey. Rescala et al. (2024) examined the reliability of LLM-as-a-judge for evaluating persuasiveness. They introduced two datasets, PoliProp and PoliIssue, featuring carefully selected debates with balanced and sufficiently long utterances. Both humans and LLMs were asked to assess the persuasiveness of arguments, and evaluate how demographic information might influence a person’s stance, and the persuasiveness of an argument. Their findings suggest that LLMs perform comparably to human judges, highlighting the promise of AI as Persuasion Judge. However, early results from Bozdag et al. (2025) caution against overreliance on this approach. Using the persuasion dataset from Durmus et al. (2024), they found that LLMs achieved only around 55% accuracy in ranking tasks, indicating a limited alignment with human judgments. These mixed outcomes suggest that while AI as Persuasion Judge is a promising direction, it still requires careful calibration and validation before it can be considered a robust substitute for human evaluation.
Dataset or Framework Relevant
Papers
Target Metric Rule‐based? LLM‐as‐a‐judge?
UKPConvArgStrict (Habernal and Gurevych, 2016), (Simpson and Gurevych, 2018), (Toledo et al., 2019) Argument Pairwise Classification
UKPConvArgRank (Habernal and Gurevych, 2016), (Simpson and Gurevych, 2018), (Toledo et al., 2019) Argument Ranking Scoring
IBMPairs (Toledo et al., 2019) Argument Pairwise Classification
IBMRank (Toledo et al., 2019) Argument Ranking Scoring
IBMRank-30k (Gretz et al., 2020) Argument Ranking Scoring (Pointwise Quality)
Persuasive-Pairs (Pauli et al., 2025) Argument and LLM Ranking Scoring
PersuasionBench (Singh et al., 2024) LLM Conventional metrics (e.g. BLEU, ROUGE, etc), LLM‐as‐a‐judge, Human Evaluation
The Persuasive Power of Large Language Models* (Breum et al., 2023) LLM LLM-as-a-judge
PMIYC (Bozdag et al., 2025) LLM Change in Agreement (LLM Self-Reported)
ChangeMyView (OpenAI, 2024b) LLM Human Evaluation
Persuasion Parallel Generation Evaluation (OpenAI, 2024b) LLM Human Evaluation
MakeMePay (OpenAI, 2024b) LLM Number of Payments
MakeMeSay (OpenAI, 2024b) LLM Game Winrate
Among Them (Idziejczak et al., 2025) LLM Game Winrate
Table 2. Datasets or frameworks for persuasiveness evaluation. For those without a formal name, the name of the paper is listed.
3.3. LLM Persuasiveness
There is growing interest in assessing the persuasive and manipulative capabilities of language models. LLMs can tailor arguments to specific audiences, reason through complex moral issues, and generate persuasive content that closely mimics human thought. This has triggered a paradigm shift in computational persuasion and reshaped the way algorithms influence human reasoning and decision-making. As a result, reliably assessing the persuasive capabilities of LLMs has become an urgent priority.
Unlike the evaluation of argument persuasiveness, where arguments are predefined by humans, LLM persuasiveness evaluation focuses on model-generated content in both single-turn and multi-turn dialogues. In this survey, we refer to this setting as AI as Persuader, where the language model acts as the source of persuasion, aiming to influence a human’s beliefs, attitudes, or behaviors through generated arguments or dialogue. Leading companies such as OpenAI (OpenAI, 2024b), Anthropic (Durmus et al., 2024), and DeepMind (Phuong et al., 2024) have investigated the persuasiveness of their models to better understand the risks associated with deploying these highly conversational systems. However, evaluating the persuasive abilities of LLMs is not straightforward, as research explores a variety of evaluation methodologies and experimental setups.
3.3.1. Human Evaluation
Understanding the dynamics of AI to Human persuasion is crucial, as most proprietary models are designed for individual users. Given this, using human subjects to assess the persuasiveness of LLMs is a natural approach. Durmus et al. (2024) assessed model persuasiveness by measuring human agreement with a claim before and after exposure to persuasive arguments generated by Claude models. This single-turn setup reveals that larger models are generally more persuasive, with deceptive techniques emerging as particularly effective. In contrast, Phuong et al. (2024) introduced a suite of multi-turn evaluation benchmarks, including Money Talks, Charm Offensive, Hidden Agenda, and Web of Lies. These involve interactive persuasion tasks such as persuading users to donate, impersonating a friend, manipulating users to take suspicious actions on the computer, and promoting false beliefs. Unlike single-turn approaches, these multi-turn evaluations are expected to capture more dynamic and complex forms of persuasion.
Similarly, Salvi et al. (2024) evaluated model persuasiveness through a debate game involving both human-human and human-LLM interactions. They found that participants were more likely to change their stance after engaging with GPT-4, regardless of whether the model had access to personal information. OpenAI (2024b) introduced two human evaluation benchmarks for assessing LLM persuasiveness. In the first setup, they used successful persuasion samples from the r/ChangeMyView subreddit as prompts, then asked LLMs to generate persuasive arguments in response. Human annotators scored both the original human arguments and the LLM-generated ones. The primary evaluation metric, referred to as the AI persuasiveness percentile relative to humans, measures the probability that a randomly chosen LLM-generated argument is rated as more persuasive than its human counterpart. The second benchmark, called Persuasion Parallel Generation, presents annotators with two arguments generated by different models for the same prompt. Annotators select the more persuasive argument, and the resulting win rates are used to compare model performance. Their results indicate that LLMs do not significantly outperform humans, and that newer models such as o1 do not show substantial gains in persuasiveness over earlier generations.
Interestingly, similar lines of research have emerged within the social sciences. Huang and Wang (2023) conducted a meta-analysis examining whether artificial intelligence is more persuasive than humans. Using statistical methods, they compared persuasion outcomes—such as changes in perception, attitudes, intentions, and behaviors—between human-AI and human-human interactions. Their findings suggested that LLMs are about as persuasive as humans. Focusing more specifically on political communication, Goldstein et al. (2024) found that AI-generated propaganda is already as persuasive as real-world examples. Similarly, Bai et al. (2025) showed that AI performs on par with humans in crafting messages aimed at shaping public attitudes toward policy issues.
However, using human subjects for the evaluation of persuasion presents several challenges, echoing the data‐collection bottleneck often highlighted in HCI research (Hämäläinen et al., 2023). First, human perception of persuasion is highly diverse: An argument that is compelling to one person may be far less persuasive to another. To account for this variability and ensure generalizability, studies must carefully plan and recruit sufficiently large and representative participants. Another major limitation is scalability. Human-subject research is inherently resource-intensive, both costly and time-consuming, making it challenging to evaluate persuasion across all conceivable scenarios and derive meaningful insights. Additionally, given the frequent release of increasingly advanced models, constantly relying on human participants for persuasion assessment is impractical. Finally, evaluating harmful persuasion with human subjects presents significant ethical challenges. Designing experiments that adhere to ethical standards and safeguard participants, particularly when dealing with manipulative, toxic, or biased content, is inherently difficult. In light of these challenges, we advocate for the development of automated systems for measuring persuasiveness—systems that offer greater scalability while also ensuring human safety.
3.3.2. Automatic Evaluation
As persuasiveness becomes increasingly recognized as a key competency of large language models, there is growing interest in scalable methods for evaluating this capability. To facilitate consistent and efficient assessment across different models, recent efforts have focused on developing reliable and comprehensive frameworks for automated persuasiveness evaluation.
Singh et al. (2024) proposed PersuasionBench and PersuasionArena to evaluate LLMs’ persuasive abilities using tweet pairs from enterprise accounts that differ in engagement. They introduced simulative tasks (e.g., predicting or comparing tweet persuasiveness) and generative transsuasion tasks, defined as transforming non-persuasive content into persuasive content under controlled conditions to improve engagement. Their evaluations spanned standard NLP metrics, human judgments, and an Oracle language model. Another work in this direction, Pauli et al. (2025), trained a regression model to score the relative persuasiveness of LLM-generated rewrites, using human-annotated text pairs collected across diverse domains.
An alternative evaluation approach frames persuasion as a conversation between two models. Breum et al. (2023) drew from social pragmatics to prompt persuader models accordingly and used binary feedback from persuadee models to indicate persuasive success. Building on this idea, Bozdag et al. (2025) introduced PMIYC, a framework for evaluating persuasiveness through multi-turn interactions between LLM-based persuaders and persuadees. In PMIYC, the persuadee provides a 1–5 score after each turn, enabling analysis of both the persuader’s effectiveness (AI as Persuader) and the persuadee’s susceptibility (AI as Persuadee) based on stance shifts. This design also facilitates evaluating how vulnerable a model is to persuasive influence. Complementing these setups with a mix of human and automated evaluation, Borah et al. (2025) assessed LLM persuasiveness across demographic dimensions, showing that both susceptibility and persuasive impact varied with demographic attributes and diversity in target audiences.
Building on the conversational persuasion evaluation paradigm, OpenAI (2024b) proposed domain-specific scenarios to assess persuasive abilities. In MakeMeSay, the persuader is given a secret codeword and must steer the persuadee into saying it; in MakeMePay, the persuader adopts the role of a con artist attempting to elicit a payment. Similarly, Zhu et al. (2025) introduced a multi-agent bargaining task, where groups of LLM sellers attempt to persuade LLM buyers to purchase a product. Idziejczak et al. (2025) explored persuasion in a social deception game inspired by Among Us, where LLMs played as either crewmates or impostors. Persuasive behavior emerged through dialogue, which was later annotated by another LLM using predefined persuasion strategies to assess how models influenced one another within the game.
While early findings suggest that LLMs can be as persuasive as humans—and in some cases more so, particularly as they improve in reasoning, personalization, and strategic communication, the underlying mechanisms driving this persuasiveness remain poorly understood. Current research has revealed that different evaluation setups may yield inconsistent or even contradictory results, with some models outperforming others in one framework while falling short in another. LLM-based evaluations may also capture persuasion between models rather than effects on humans, risking circularity. This highlights the complex, multidimensional nature of persuasive ability and raises important concerns about the increasing influence of ”smarter” models on beliefs and behaviors. As a result, developing a unified and trustworthy evaluation framework that integrates diverse test cases and assesses persuasive skills across multiple dimensions is a pressing research challenge. Such a framework would not only support more self-consistent and interpretable evaluations but also help explain why certain models excel in specific contexts while lacking in others, ultimately informing safer and more responsible deployment of persuasive AI systems.
4. Generating Persuasion
AI-driven systems designed to generate persuasive content are becoming increasingly powerful and widespread. Persuasion can play a pivotal role in various domains, including advertisements, healthcare promotion, recommendation systems, political campaigns, and more. As organizations and individuals seek to increase their influence, the need for generating persuasive content has grown significantly, especially with the advent of LLMs. This section explores prior work on enhancing persuasive capabilities, examines key influencing factors, and highlights applications of persuasion in settings like negotiation and debate.
4.1. Methods
Section 2.2.1 introduced taxonomies of persuasive strategies that LLMs can adopt to enhance the effectiveness of their outputs. Building on this foundation, recent work has focused on developing methods to improve LLMs’ ability to generate persuasive content. In this subsection, we review key approaches proposed to enhance the persuasive capabilities of LLMs, including prompting techniques, incorporation of external knowledge, fine-tuning, and reinforcement learning (See Table 3). Beyond neural approaches, structured reasoning methods from computational argumentation offer an alternative way to model persuasive generation. Argumentation-based chatbots have proven particularly effective in persuasion, where explicit reasoning and dialogue structure support goal-directed interaction (Castagna et al., 2024). Integrating such methods with LLMs could enable hybrid systems that combine the fluency of neural models with the interpretability and deliberative rigor of formal argumentation.
Method Work Persuasive Factor Dataset Brief Description
Prompting (Pauli et al., 2025) Assigns persona to LM
(Durmus et al., 2024) Instructs the use of specific persuasive strategies
(Zeng et al., 2024) Adversarial intensity Instructs the use of specific persuasive strategies
(Xu et al., 2024a) Misinformation Instructs the use of specific persuasive strategies
Incorporating External Information (Salvi et al., 2024) Personalization Personal information of participants
(Zhang and Zhou, 2025) DailyPersuasion (Jin et al., 2024) Observations for persuadee’s mental state
(Tiwari et al., 2022a) Personalization PPMD (Tiwari et al., 2022a) Selects a persuasive strategy based on the user profile
(Tiwari et al., 2023) Personalization PPD (Tiwari et al., 2023) Selects a persuasive strategy based on the user profile
(Cima et al., 2024) Personalization Reddit User comment history and summary of user’s style
(Furumai et al., 2024) Factuality Fact-checking and retrieval for persuasion
(Karande et al., 2024) Factuality Strategy planning and information retrieval
Finetuning (Chen et al., 2022) Factuality, Engagement PersuasionForGood (Wang et al., 2019) Finetunes BART for conditional persuasion generation
(Liu et al., 2021) Empathy ESConv (Liu et al., 2021) Finetunes models for emotional support generation
(Jin et al., 2024) Personalization DailyPersuasion (Jin et al., 2024) Finetunes Llama-2-Chat to reason for user intentions
(Singh et al., 2024) Tweets Instruction-finetunes Vicuna-1.5 13B
Reinforcement Learning (Shi et al., 2021) PersuasionForGood (Wang et al., 2019) Trains persuasive models in an RL setting
(Samad et al., 2022) Empathy PersuasionForGood (Wang et al., 2019) Trains empathetic persuasive models in an RL setting
(Tiwari et al., 2022b) Personalization DevPVA (Tiwari et al., 2022b) Rewards persuasion based on persona of the user
(Mishra et al., 2022) Empathy PersuasionForGood (Wang et al., 2019) Trains empathetic persuasive models in an RL setting
(Hong et al., 2025) ESConv (Liu et al., 2021), PersuasionForGood (Wang et al., 2019) Offline RL training for increased persuasion
Table 3. Overview of persuasion generation methods. Persuasive Factor refers to any specific aspect of persuasiveness (e.g., personalization, empathy, factuality) that the work aims to modify to enhance persuasive impact (– means no specific factor was targeted).
4.1.1. Prompting
One foundational method for enhancing persuasive generation is prompt engineering (Rogiers et al., 2024). Even a simple instruction to rewrite text more persuasively can lead to substantial improvements in persuasiveness (Pauli et al., 2025). Notably, persuasion can also be influenced indirectly through prompting; assigning a particular persona to an LLM, even without explicitly referencing persuasion, has been shown to increase the persuasiveness of its responses (Pauli et al., 2025) .
Another approach is to prompt LLMs with specific persuasive strategies. Durmus et al. (2024) examined how different types of prompts affect the persuasiveness of model outputs by experimenting with four prompting strategies: (1) prompting the model to construct compelling arguments, (2) adopting the role of an expert and using rhetorical techniques such as pathos, logos, and ethos, (3) encouraging logical reasoning, and (4) permitting deception, including fabricating evidence, to boost persuasive impact. Their findings showed that while effectiveness varies across model architectures, prompts that encourage logical reasoning or allow for deception tend to produce more persuasive arguments overall. Along similar lines, Zeng et al. (2024) designed a comprehensive taxonomy of persuasion strategies, then used these strategies to generate persuasive adversarial prompts which proved to be remarkably effective in jailbreaking LLMs. Using persuasive strategies has also demonstrated effectiveness in misinformation propagation, where persuasive strategies effectively manipulate LLMs into endorsing false information (Xu et al., 2024a).
Prompting has also proven effective in multi-agent systems for generating persuasive interactions. Several studies have relied solely on prompting to explore the persuasive capabilities of LLMs in simulated multi-agent dialogues (Bozdag et al., 2025; Breum et al., 2023). Beyond evaluation, prompting is increasingly used as a tool for scalable data generation. For instance, Ma et al. (2025) introduced a multi-agent prompting framework in which LLMs assume specialized roles (e.g., persuader, quality monitor, annotator) to simulate structured persuasive dialogues. Their method integrates layered quality checks and strategy-aware prompting to construct large-scale, high-quality persuasion datasets with minimal human supervision. However, the use of prompting to elicit persuasive reasoning, such as chain of thought prompting or structured argumentative strategies, remains underexplored.
4.1.2. Incorporating External Information
While prompting offers a good baseline for persuasive generation, incorporating external information can substantially enhance persuasiveness by introducing additional dimensions such as personalization and factual grounding. Models can produce more tailored and credible persuasive content by leveraging contextual knowledge, including user profiles, background information, or external evidence. Effective persuasion involves not only implementing persuasive strategies but also selecting techniques that align with the characteristics of the recipient. Lukin et al. (2017) and Wang et al. (2019) showed that accounting for personality traits is critical, as individuals with different personality types tend to respond differently to specific persuasive strategies. Building on this insight, Jin et al. (2024) proposed a method where LLMs begin by identifying and summarizing the user’s intent before reasoning about the most suitable strategy to apply. This process enables a more adaptive and responsive persuasive system.
Incorporating personal information about the user extends beyond just strategy selection. Personalization based on users’ psychological traits, such as personality, political ideology, moral foundations, and behavioral patterns, has been shown to significantly enhance persuasive impact (Matz et al., 2024; Kaptein et al., 2015). This approach is effective across various domains, including targeted advertisements and political messaging around climate action (Matz et al., 2024; Simchon et al., 2024; Meguellati et al., 2024). Salvi et al. (2024) showed that providing LLMs with personal information during human interaction increases their persuasiveness. Research by Ruiz-Dolz et al. (2024) demonstrated that user-specific persuasive policies significantly improve the persuasiveness of argument-based systems. Zhang and Zhou (2025) emphasized the importance of considering the user’s mental state when constructing persuasive arguments. Tiwari et al. (Tiwari et al., 2022a, 2023) showed that combining dialogue context with user profile information provides further persuasive benefits. In parallel, other research highlighted the value of counterarguments that directly engage with the user’s concerns or beliefs (Cima et al., 2024; Hunter et al., 2019).
Beyond personalization, grounding persuasive content in factual information is essential for building credibility and enhancing persuasive impact. Incorporating verified information into conversations not only strengthens an agent’s arguments but also improves the overall user experience (Chen et al., 2022). PersuaBot (Furumai et al., 2024) addresses factuality by generating an initial response with an LLM, decomposing it into segments, and verifying each segment using information retrieval techniques. This pipeline mitigates hallucinations and reduces unsupported claims, thereby increasing both persuasiveness and factual accuracy. Similarly, comprehensive multi-agent systems such as the one proposed by Karande et al. (2024) incorporate external knowledge through modular components that cover retrieval, response analysis, strategy selection, and fact validation, to construct more robust and credible persuasive agents.
4.1.3. Finetuning
Finetuning LLMs on persuasive datasets offers a powerful way to enhance their persuasive capabilities beyond what prompting alone can achieve. This approach allows models to internalize persuasive patterns and strategies at a more fundamental level. In the negotiation domain, Lewis et al. (2017) trained a model that is based on four recurrent neural networks on negotiation datasets, and demonstrated significant improvements in negotiation ability. Extending finetuning to emotional support contexts, Liu et al. (2021) introduced the Emotion Support Conversation dataset (ESConv) and showed that finetuning BlenderBot (Roller et al., 2021) and DialoGPT (Zhang et al., 2020) on this data enhances models’ ability to provide empathetic responses and sustain engagement.
Chen et al. (2022) finetuned BART-large (Lewis et al., 2020) on the PersuasionForGood dataset (Wang et al., 2019), achieving significant gains in persuasive quality. More recently, Jin et al. (2024) developed PersuGPT by training LLaMa-2-13B-Chat on the DailyPersuasion dataset, which includes rich reasoning traces. Combining conventional finetuning with Direct Preference Optimization (DPO) (Rafailov et al., 2023), their model learns to infer user intent, select appropriate persuasive strategies, and anticipate future user responses and rewards. Singh et al. (2024) proposed an instruction-based finetuning approach to enhance LLM persuasiveness. They finetuned Vicuna-1.5 13B to rewrite inputs to be more persuasive and showed that the resulting model outperforms the base version and can rival much larger LLMs in persuasiveness. Finetuning presents strong potential for developing more persuasive language models, yet remains relatively underexplored. With more carefully curated training data and domain-specific finetuning, future work can yield models that are better aligned with persuasive goals across a range of applications, including healthcare, education, and advertising.
4.1.4. Reinforcement Learning
Reinforcement learning (RL) approaches are also highly effective for developing persuasive models, enabling more nuanced control over multiple persuasive dimensions through carefully designed reward functions. Several studies have shown the effectiveness of RL in enhancing persuasive qualities. Samad et al. (2022) augmented the PersuasionForGood dataset with emotion annotations and trained an agent using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm for more empathetically engaging and persuasive responses. Their reward function combined penalties for repetition and topical drift with incentives for emotional engagement and appropriate persuasive strategy use. Similarly, Mishra et al. (2022) refined emotional expression and politeness in model persuasive generations. They employed PPO with a reward that incorporates signals from persuasion, emotion, politeness strategy consistency, dialogue-coherence, and non-repetitiveness.
Tiwari et al. (2022b) applied RL to task-oriented dialogue agents, showing that persona-aware rewards could improve persuasive effectiveness toward shared goals. Similarly, Shi et al. (2021) also trained a GPT-2-based persuader with DialGAIL, which uses RL to address the issues of repetition, inconsistency, and task relevance. Most recently, (Hong et al., 2025) introduced an approach leveraging hindsight regeneration for training persuasive dialogue systems. They first simulate human-agent conversations and then retrospectively optimize agent responses based on the success of conversation outcomes, which is called hindsight regeneration. Finally, they employ offline RL on this data, resulting in improved persuasive performance. Although RL provides a powerful framework for shaping persuasive behavior, it remains underexplored. Future research should prioritize the development of more expressive and reliable reward functions that can capture subtle aspects of persuasive quality, such as ethical alignment, user receptivity, and long-term impact. In addition, there is substantial room to apply RL-based persuasion in many other domains. Methods like RLHF and RLAIF offer new opportunities for aligning persuasive generation with human values and feedback.
4.2. Applications of Persuasion
Persuasive LLMs hold promise across a wide range of domains where shaping beliefs or influencing decisions is essential, including areas like advertising, education, mental health counseling, and political communication. Many of these applications remain underexplored, and differ widely in their goals, audiences, and ethical implications. In this section, we focus on three key settings that highlight distinct dynamics of persuasion with LLMs: negotiation, which exemplifies cooperative persuasion; debate, which involves adversarial reasoning; and jailbreaking, which reveals how persuasion can be used to exploit model vulnerabilities. These cases illustrate not only the potential of persuasive LLMs but also the challenges in aligning them with intended outcomes.
4.2.1. Negotiation
Negotiation is a semi-cooperative setting, where intelligent agents with different goals try to reach a mutually acceptable solution (Lewis et al., 2017). Persuasion plays a central role in negotiation, where agents must influence their counterparts’ preferences, beliefs, or concessions to reach mutually beneficial agreements. In both agent-agent and agent-human settings, persuasive strategies are employed to build rapport, justify proposals, and strategically guide the dialogue toward favorable outcomes. In their work, Lewis et al. (2017) presented a conversational dataset where humans negotiate on a multi-issue bargaining task and try to reach an agreement. They demonstrated that training dialogue agents with reinforcement learning, where the reward is based on the negotiation outcome, is more effective than supervised methods for improving negotiation abilities. Additionally, they introduced a form of planning for dialogue called dialogue rollouts, where an agent simulates dialogues during decoding to estimate the reward of utterances. They found that their agents demonstrate sophisticated negotiation strategies and also learn to be deceptive. In their work, Keizer et al. (2017) compared five different conversational agents that negotiate trades with humans in an online version of the game ”Settlers of Catan”. The different agents employ different negotiation strategies and demonstrate that persuasion leads to improved win rates. They also demonstrate that strategy selection based on deep reinforcement learning is effective as well.
Similarly, He et al. (2018) observed two agents that interact with each other to bargain on goods. They proposed an approach based on coarse dialogue acts, where they separate strategy and generation. They set the strategy using supervised learning, reinforcement learning, or domain-specific knowledge. This is used to select a dialogue act, then they generate a response. They show that this proposed method demonstrates more human-like negotiation and a higher task success rate based on human evaluations. Later works bring more focus to planning and persuasive strategies for negotiation. DialoGraph (Joshi et al., 2021) incorporated Graph Neural Networks to model sequences of strategies for negotiation, and incorporate them into a negotiation dialog system. (Chawla et al., 2021a) presented a dataset of negotiation dialogues and a multi-task framework for recognizing the persuasion strategies used in an utterance. Most recently, Bianchi et al. (2024) created NegotiationArena, a framework used to assess how well LLMs negotiate with each other to maximize their profits and resources. The results showed that GPT-4 is overall the best negotiating LLM. Their experiments highlight some interesting findings, such as cunning and desperate behaviors increase win rate and payoff.
The increasing persuasive capabilities of LLMs open new possibilities for negotiation, where models can serve as collaborative agents, mediators, or decision-support tools that guide discussions toward favorable outcomes. However, these enhanced abilities also raise questions about how susceptible such systems may be when operating in multi-agent environments or interacting with humans who attempt to manipulate them. Understanding both the strengths and vulnerabilities of persuasive LLMs will be essential for building negotiation agents that are both effective and robust.
4.2.2. Debate
In contrast to the cooperative dynamics of negotiation, debate is inherently adversarial, with agents taking opposing positions and attempting to persuade an audience or judge of their stance. A foundational step toward computational debate was Slonim et al. (2021)’s Project Debater, an autonomous system capable of mining, organizing, and rebutting arguments in live debates against human experts, illustrating that large-scale argument generation and persuasion can be implemented in AI. In subsequent research, Michael et al. (2023) explored prompting two expert LLMs to debate answers to reading comprehension questions to persuade a non-expert human judge. The back-and-forth exchange exposes flaws and inconsistencies in each model’s reasoning, ultimately helping the human identify the correct answer. This work highlights the potential of debate as a supervision mechanism for aligning expert models and underscores the importance of studying persuasion as a core component of such interactions. Du et al. (2024) proposed a multi-agent debate framework in which LLMs iteratively propose and critique responses across several rounds before converging on a final answer. This debate-driven strategy was shown to improve mathematical reasoning, strategic thinking, and the factual correctness of generated content. Together, these studies demonstrate that debate not only facilitates persuasion but also contributes to model alignment and output reliability through adversarial reasoning.
In debate settings, persuasive LLMs offer a powerful mechanism for strong, quality argumentation that can use complex reasoning. As these systems become more adept at argumentation, their susceptibility to persuasion poses risks, especially in adversarial debates where subtle rhetorical manipulation may distort outcomes. Future work must address how to maintain argumentative rigor and guard against manipulation in AI-driven debate systems.
4.2.3. Jailbreaking
Beyond understanding the inherent persuasive capabilities of LLMs, it is crucial to look at how strong persuasive techniques–used by either humans or LLMs–can be strategically employed to exploit vulnerabilities in models. This scenario is viewed as ”AI as Persuadee”, in which target models can be swayed by carefully crafted persuasive prompts. Such prompts can jailbreak a target LLM into bypassing established safety mechanisms and leading to the generation of harmful, unsafe, or incorrect content. This section reviews recent studies on how persuasion-based techniques can jailbreak LLMs, highlighting the associated risks and potential mitigation.
Harmful/Toxic Content Generation. Jailbreaking LLMs means subverting their safety measures and producing harmful, biased, or sensitive generations. Recent work (Singh et al., 2023; Zeng et al., 2024) highlight a vulnerability in LLMs, where persuasion-based jailbreaks can bypass existing safety mechanisms. Singh et al. (2023) demonstrates that LLMs can be deceived using tailored persuasion strategies, where persuasive prompts are systematically crafted by leveraging authority, trust, and social proof. Zeng et al. (2024) develop a Persuasive Paraphraser which rephrases harmful queries using a persuasion taxonomy grounded in social science. A key finding is that larger LLMs exhibit greater susceptibility to such persuasive attacks, likely due to their improved contextual reasoning and user engagement capabilities. They also evaluate existing defense mechanisms, revealing that existing mutation- and detection-based defenses are insufficient, while adaptive system-level defenses (e.g., reinforcement against persuasion) provide some mitigation, but remain limited. Moreover, Li et al. (2024) introduces multi-turn prompts with the persuasion technique, where users repeatedly apply persuasion techniques over conversations, leading LLMs to generate harmful content in multi-turn dialogues. These findings emphasize the need to rethink AI safety in the context of human-like communication, advocating for more robust defenses against social science techniques to prevent unintended model behavior.
Misinformation Generation. Xu et al. (2024a) examine the vulnerability of LLMs to well-crafted persuasive prompts. Their study employs a multi-turn dialogue framework to demonstrate that iterative application of persuasive strategies–such as repetition and nuanced rhetorical appeals–can effectively jailbreak LLMs, leading to significant shifts in their factual responses but containing misinformation. Moreover, (Bozdag et al., 2025) introduces a framework for evaluating the persuasive effectiveness and susceptibility of LLMs over multi-turn conversation interactions, showing that LLMs generally become more susceptible in multi-turn conversation than in single-turn conversation. In essence, their work underscores how sequentially deployed persuasive prompts gradually erode the robustness of LLMs, revealing critical weaknesses in current safeguarding mechanisms against misinformation generation.
5. Safeguarding Persuasion
While persuasive communication can be beneficial, such as helping users make informed, healthy, or prosocial choices, it also carries risks of manipulation, bias, and erosion of autonomy. Responsible AI frameworks emphasize transparency, accountability, fairness, and non-maleficence (Radanliev, 2025; Radanliev et al., 2024). These principles extend naturally to persuasive models, which must balance efficacy with respect for user consent and epistemic integrity. A central issue in persuasion ethics is distinguishing virtuous vs. vicious persuasion. As argued by Godber and Origgi (2023), legitimate (virtuous) persuasion engages people’s deliberative and epistemic capacities, enabling reflective choice and preserving intellectual autonomy. Manipulative or propagandistic (vicious) persuasion instead exploits cognitive or emotional vulnerabilities to advance the persuader’s interests at the expense of the audience’s welfare. From a computational perspective, Bassi et al. (2024) observe that current NLP and ML systems seldom encode a principled distinction between ethical and unethical persuasion. Embedding such conceptual clarity into computational frameworks is essential for developing persuasion systems accountable to humanistic values.
As persuasive AI systems become increasingly influential, it becomes critical and urgent to ensure their responsible deployment. Safeguarding persuasion involves a wide range of challenges, including but not limited to understanding when and how persuasion should be employed, identifying and mitigating unwarranted or manipulative persuasive tactics, and calibrating the susceptibility of AI systems to persuasion without making them overly rigid. In this section, we review early work aimed at reducing the risks of unsafe persuasion and developing models that can more selectively decide when to accept, reject, or generate persuasive content. Compared to the other two areas covered in this survey—Evaluating Persuasion (§ 3) and Generating Persuasion (§ 4)—although equally important, this area has the least amount of prior work, in part because concerns around the safe use of persuasive AI have only recently begun to take shape with the advance of increasingly persuasive models.
5.1. Mitigating Unsafe Persuasion
Persuasion is not inherently harmful. In fact, effective AI persuaders can be beneficial in many contexts. For example, it can power positive change, nudging users toward healthier habits or aligning multi‐agent systems with optimal decisions. Yet as persuasive power grows, so too does the risk of misuse. A model tailored to individual preferences can subtly erode autonomy, reinforcing biases or nudging users toward risky or illegal behavior. For instance, Burtell and Woodside (2023) warned that hyper-personalized dialogue may circumvent traditional decision-making safeguards, while El-Sayed et al. (2024) proposed methods for persuasion harmfulness mitigation, such as red-teaming.
Persuasion risks underscore the need for both technical and regulatory solutions. Currently, there are few systems in place to ensure that persuasive AI models are used safely and responsibly. The first step toward mitigating unsafe persuasion is the ability to detect when persuasion is taking place, as discussed in § 3.1. But more importantly, systems must learn to distinguish between generating persuasive content that is helpful or ethical and content that is manipulative, coercive, or otherwise harmful. Addressing this challenge requires a combination of approaches: model interpretability can help reveal the reasoning behind persuasive outputs, content filtering can prevent harmful language or tactics from being generated, and ethical frameworks or regulatory guidelines can establish clear boundaries around what constitutes acceptable persuasion. Together, these components support the development of persuasive systems that are transparent, accountable, and aligned with societal values.
5.2. Selective Acceptance of Persuasion
In addition to determining when and how to persuade, models must also develop the capacity to discern when to accept influence and when to resist it. We refer to this capability as selective acceptance of persuasion–the ability of a model to intelligently evaluate incoming persuasive attempts and decide whether to adopt or reject them based on context, alignment, and reliability. Dutt et al. (2021) introduced a novel line of work that shifts focus to resistance strategies, which are the methods individuals use to counter or reject persuasive attempts. Their framework, ResPer, drew from cognitive and social psychology to identify and categorize these strategies in real-time dialogue. Their approach revealed how resistance unfolds dynamically within conversations, exposing asymmetries in power and influence. Crucially, their findings showed that resistance strategies are often more predictive of conversation outcomes than persuasion tactics themselves. Such insights could inform the design of AI systems that are more discerning in responding to persuasive input, making them more receptive to beneficial influence while remaining resilient to manipulative attempts.
Stengel-Eskin et al. (2025) focused on when models should do so by proposing Persuasion-Balanced Training (PBT), a method that teaches models to selectively accept helpful (positive) persuasion while resisting harmful (negative) attempts. They simulate dialogues between persuader and persuadee agents, where the goal is to reach a consensus on a given question. These interactions are structured as recursive dialogue trees, from which preference data is derived. Using DPO loss (Rafailov et al., 2023), they train models to favor dialogue paths that lead to correct answers, regardless of whether persuasion was accepted or resisted. Their findings show that PBT significantly reduces misinformation and flipflopping, and improves collaboration in multi-agent debates, outperforming models trained only to resist or accept persuasion. Taken together, ResPer and PBT offer complementary perspectives on selective acceptance: the former provides a taxonomy of resistance strategies, while the latter offers a learning framework for context-sensitive influence evaluation.
Still, real-world deployments pose additional challenges. In open-ended interactions with unfamiliar agents, intent is hard to infer, and deciding to accept or reject persuasion is inherently more nuanced. Sycophancy, the tendency to favor flattering or agreeable responses, presents an added challenge to learning selective persuasion. Both humans and preference models have been shown to favor convincingly written sycophantic responses over those that correct false beliefs in many cases (Sharma et al., 2024). This could lead to models falsely accepting harmful persuasion while still rejecting helpful ones. To build trust in persuasive systems, we must deepen our understanding of what models internalize during training and how they make influence-sensitive decisions. Red-teaming has emerged as a powerful tool for identifying such vulnerabilities. Scalable LLM-to-LLM red-teaming frameworks enable adversarial agents to systematically provoke undesirable behavior in target models before deployment (Perez et al., 2022). In persuasion-specific settings, red-team agents can manipulate models into accepting misleading influence, generating examples of where discernment fails. These adversarial conversations can then be used to train more robust models, helping them recognize patterns that lead to harmful persuasion and improve their ability to respond appropriately in real-world interactions.
6. Persuasion Beyond English Text
While our survey primarily focuses on persuasion in English text using large language models, much of the global diversity in persuasive communication lies beyond this scope. Persuasion manifests differently across modalities, languages, and cultural contexts, each with its norms, strategies, and challenges. In this section, we highlight pioneering research that explores these broader dimensions.
Mutimodal Persuasion. Although most computational research on persuasion has focused on textual data, human persuasion is inherently multimodal and often conveyed through speech, visuals, or a combination of modalities. Understanding how persuasive cues operate across these forms is crucial for developing more generalizable and context-aware persuasive systems. Early work by Park et al. (2014) explored multimodal persuasion by predicting persuasiveness in movie review videos using speech, text, and visual features. Building on this, Bai et al. (2023) proposed the Multimodal Persuasion Prediction (M2P2) framework, which models both the outcome and intensity of persuasion in debates. Later, Liu et al. (2022) introduced the ImageArg dataset, designed to support image-based persuasion detection. Other efforts have leveraged transcribed video data from social deduction games to study persuasive behaviors in complex, interactive settings (Lai et al., 2023). Researchers have also explored persuasive strategies in visual advertisements (Kumar et al., 2023), and in meme-based content through shared tasks organized at SemEval (Dimitrov et al., 2021, 2024). Xu et al. (2024b) demonstrated that vision-language models can be poisoned to generate misleading yet persuasive multimodal narratives. These efforts lay the groundwork for future research in persuasive content generation and safety mechanisms for multimodal persuasion.
Multilingual and Culture-Aware Persuasion. As with many NLP tasks, research on persuasion has predominantly focused on English, thanks to the availability of large-scale datasets and language resources. However, real-world persuasive communication spans a multitude of languages and cultural contexts, where data scarcity and linguistic diversity introduce additional challenges. The task of modeling persuasion becomes significantly more complex in multilingual and cross-cultural settings, hence, it remains relatively underexplored. One of the earliest efforts toward multilingual persuasion modeling is presented by pöyhönen2022multilingualpersuasiondetectionvideo, who introduced a dataset of persuasion in video content across multiple languages. More recently, Piskorski et al. (2023) organized a SemEval shared task on persuasive strategy classification in a multilingual news corpus. Substantial research gaps remain, and fundamental tasks such as persuasion detection and strategy classification have yet to be fully extended to non-English languages. Furthermore, it remains an open question how cultural and linguistic context influences the choice and effectiveness of persuasive strategies. Addressing these gaps is essential for building persuasive systems that are culturally sensitive and globally applicable.
7. Challenges & Future Directions
In this survey, we have reviewed prior research in computational persuasion, focusing on three key perspectives: AI as Persuader, AI as Persuadee, and AI as Persuasion Judge. While significant progress has been made in understanding and modeling various aspects of persuasion, numerous open challenges and promising directions remain. The rapid development and widespread adoption of large language models offer new opportunities while also raising urgent questions, necessitating a reexamination of core assumptions and a rethinking of methodologies in persuasion research.
7.1. AI as Persuader
The generation of persuasive content using LLMs offers significant promise, but also presents serious risks. Although recent studies have shown that LLMs are becoming increasingly effective at persuasion (Durmus et al., 2024; OpenAI, 2024b, 2025; Bozdag et al., 2025; Singh et al., 2024), current research has merely scratched the surface, leaving much to be explored in understanding the capabilities, limitations, and broader implications of persuasive LLMs.
7.1.1. Evaluation of Persuasiveness
In this survey, we have described a range of evaluation frameworks aimed at assessing the persuasive abilities of language models (see § 3). However, there is still a need for:
(1) Unified and Comprehensive Evaluation: As the development and release of new models accelerate, establishing standardized evaluation benchmarks that allow for rigorous and transparent comparisons becomes increasingly important. Such benchmarks would not only advance scientific understanding of persuasive capabilities but would also encourage accountability from model developers. Rather than relying on proprietary or self-defined risk categories, shared evaluation frameworks and benchmarks would help align the community around a common understanding of the potential risks and benefits associated with persuasive LLMs. These evaluations should ideally span multiple dimensions of persuasion, such as the use of logical, emotional, or deceptive strategies (Durmus et al., 2024), conversational dynamics (both AI to Human and AI to AI) (Bozdag et al., 2025), and commercial or marketing-oriented persuasion (Wu et al., 2025), which prior work studied in isolation, leading to inconsistent findings.
(2) Automating Human-as-Subject Evaluations: Current methods for assessing persuasiveness in human-as-subject interactions rely heavily on human participants, making them costly, risky, and difficult to scale. We propose automating these evaluations through the use of well-defined user simulators that represent a diverse range of user profiles, cognitive states, and goal structures, ideally grounded in behavioral theory and real-world data. By modeling how different users might respond to persuasive content, these simulators can serve as reliable proxies for large-scale evaluations, while mitigating the risks of exposing actual humans to potentially manipulative or adversarial prompts.
7.1.2. Mechanics of Persuasion Generation
While LLMs are demonstrating notable improvements in persuasive capability (OpenAI, 2025, 2024b; Durmus et al., 2024; Phuong et al., 2024), we still lack a clear understanding of what makes them effective (or ineffective) persuaders. Much of the underlying mechanics that drive LLMs’ persuasive success remains opaque, limiting our ability to reliably predict, enhance, or safeguard their behavior. Four major research gaps are:
(1) Understanding Emergent Persuasive Behavior: When not restricted by predefined strategies, it remains unclear what kinds of persuasive techniques LLMs spontaneously adopt. Do models adapt their strategies based on the context, topic, or perceived audience profile? Are they more convincing when advocating for positions they internally “prefer,” or is persuasion success independent of any model-internal alignment? These fundamental questions about how and when LLMs persuade remain largely unexplored. Prior studies have relied on limited data (Carrasco-Farre, 2024) or primarily focused on human-designed categories of persuasion (e.g., ethos, pathos, logos), but there has been little empirical analysis of what emergent patterns LLMs exhibit when tasked with open-ended persuasion . Answering these questions is critical for building more controllable and interpretable persuasive agents.
(2) Pro-social Applications of Persuasion: After early work such as Wang et al. (2019) which examined persuasion for donation, research has rarely explored how persuasive capabilities could be leveraged for socially beneficial domains. Fields such as education, healthcare, and personal well-being remain underexamined, despite their clear societal importance. Understanding how LLMs can adapt persuasive strategies to support learning, promote healthy behaviors, or enhance mental well-being presents a critical opportunity for future work. Moreover, studying persuasion in these contexts can inform the development of safeguards that prevent coercion and promote genuinely pro-social influence.
(3) Target-Specific and Adaptive Persuasion: While LLMs show some contextual adaptation, their ability to tailor persuasive strategies to individual users remains largely unexplored. Advancing this line of research will require controlled studies where models are given user background information and tasked with generating personalized persuasive messages, with effectiveness compared against generic outputs. Future work should also explore adaptive persuasion in multi-turn settings, where models refine strategies based on user feedback or environmental signals. Earlier decision-theoretic approaches have treated persuasion as an interaction between agents with separate objectives, where user preferences are learned to guide adaptive argument choices (Donadello et al., 2022). Framed as a preference learning and alignment problem, adaptive persuasion requires models to infer and respect user values, goals, and receptivity to generate effective and ethically grounded influence.
(4) Learning Persuasive Policies through Reinforcement: It remains an open question whether LLMs can be systematically trained to improve their persuasive decision-making. Current approaches based on prompt engineering or supervised fine-tuning do not equip models with dynamic, adaptive policies for persuasion. Future research should pursue persuasive policy learning, where a policy represents the model’s learned mapping from dialogue states to persuasive actions that optimize a defined objective—be it ethical influence, pro-social encouragement, commercial engagement, or adversarial robustness. In this context, models learn to generate persuasive content through reinforcement signals that reward both efficacy and ethical responsibility. This requires designing reward models that balance persuasive success with user well-being, fairness, and trustworthiness.
7.1.3. Long Context Persuasion
Persuasion generation has primarily been studied in the context of stand-alone arguments, single-turn, or short multi-turn conversations. However, there is significant potential in developing systems capable of engaging in long-term, multi-session persuasion that is more subtle, context-aware, and reflective of natural human interactions. Such capabilities promise more natural, context-aware interactions, but also introduce new risks. Understanding how models maintain persuasive strategies across extended dialogues is critical for both advancing the field and ensuring responsible deployment. Key challenges include:
(1) Benchmarking Long-Horizon Persuasion: Existing evaluation methods focus on short interactions, not assessing persuasive success over extended conversations. Future benchmarks should track how persuasive influence evolves across multiple turns and sessions, considering factors like memory, trust dynamics, and user resistance over time.
(2) Persuasive Planning and State Tracking: Long-term persuasion requires more than turn-by-turn skill. Models must plan across dialogue horizons, adapt strategies dynamically, and maintain memory of user preferences and prior interactions. Future work should explore methods like hierarchical planning, user modeling, and memory-augmented architectures to support sustained and ethical persuasion.
7.2. AI as Persuadee
While recent work has shown that LLMs can be susceptible to persuasive prompts and adversarial strategies (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025), the boundaries and underlying mechanisms of these vulnerabilities are not well understood. Insights from steering vectors may further clarify such mechanisms (Dunefsky and Cohan, 2025), and several important directions merit further investigation.
7.2.1. Model Susceptibility to Persuasion.
It is not yet clear whether LLMs respond to persuasive strategies in ways analogous to humans. Existing studies have primarily focused on short, single-turn interactions, leaving many open questions about how models internalize influence, especially over longer contexts. Understanding model susceptibility is crucial for both advancing robustness and anticipating new forms of adversarial manipulation. The primary concerns are:
(1) Susceptibility in Long Interactions: Most existing work examines persuasion in isolated or a limited number of exchanges, but real-world attacks often unfold across extended dialogues. Research should investigate how LLMs respond when persuasive strategies are diffused over many turns, especially when combined with techniques such as many-shot jailbreaking (MSJ) (Anil et al., 2024). Studying subtle, multi-turn persuasion could reveal deeper model vulnerabilities that are not apparent in short-form settings.
(2) Effective Modalities of Agent Persuasion: While natural language is the dominant input modality, it is not yet known whether alternatives such as structured metadata, tool outputs, or interleaved code could more effectively persuade models. Exploring multimodal input may uncover novel vectors of influence and corresponding risks.
(3) Origins of Persuasive Vulnerability: It is not yet known which stages of model development (e.g., instruction tuning, reinforcement learning, alignment techniques) contribute most to susceptibility. Future work should aim to isolate these factors to better understand and ultimately protect model responses against manipulative inputs. This behavior might also be examined through models’ self-explanations of why they were persuaded, offering insight into which aspects of a persuasive argument influenced the response.
7.2.2. Selective Acceptance and Resistance to Persuasion.
Understanding model susceptibility is only the first step; equally important is developing models that can selectively accept or resist persuasion without degrading their general capabilities or instruction-following performance (Stengel-Eskin et al., 2025). Building such systems will require new training objectives, fine-tuning strategies, and evaluation protocols explicitly targeting persuasive resilience. Remaining challenges are:
(1) Defining Selectiveness in Persuasion: Persuasion resistance should not emerge from general stubbornness or degraded instruction-following, but from correctly identifying problematic persuasive attempts. We must establish clear definitions and metrics for selective robustness that distinguish between appropriate and inappropriate influence.
(2) Training Selective Persuadees: Achieving selective behavior will require novel training techniques, such as fine-tuning on adversarially generated persuasive attacks, training with annotated persuasive feedback, or using preference learning to steer model responses toward robustness while preserving flexibility.
7.2.3. Agent-to-Agent Systems: AI as Persuadee & Persuader
Research in computational persuasion has primarily examined human-AI interactions, but as multi-agent systems become increasingly common, novel safety concerns, including agent-to-agent persuasion, are likely to emerge (Hammond et al., 2025). Studying persuasion in multi-agent contexts is therefore critical. The main challenges are:
(1) Persuasion Effects in Multi-Agent Systems: Future work must examine how persuasion among agents can influence outcomes in debates, negotiations, and collaborative tasks. It is important to identify when and how more persuasive or more susceptible agents shift system behavior toward undesirable states. Moreover, it remains unclear whether agent-to-agent persuasion will resemble human-like rhetorical strategies through natural language or evolve through novel modalities and mechanisms.
(2) Inter-Agent Dynamics & Learning Profiles: Open questions include whether stronger models (e.g., larger or better-aligned ones) more easily persuade weaker ones, whether weaker agents can effectively persuade stronger ones, and how agents might adapt their trust in others through repeated interaction. Designing safeguards will require careful modeling of agent profiles, trust calibration, and learning dynamics in multi-agent ecosystems.
7.3. AI as Persuasion Judge
Aside from being a Persuader or Persuadee agent, language models also hold great potential to serve as Persuasion Judges where they detect, mitigate, and inform about persuasive attempts, recognize and differentiate harmful persuasion from beneficial encouragements.
7.3.1. Detecting & Evaluating Persuasion
Automatically identifying patterns of persuasion—such as rhetorical strategies, emotional appeals, or attempts at undue influence—is a critical use case for both classifiers and generative language models. As reviewed in § 3, prior work has explored modeling persuasion and distinguishing successful from unsuccessful persuasive attempts. However, these systems often lack alignment with human preferences or generalize poorly beyond the specific domains they were trained on. Improving detection and evaluation models remains a challenge, with several open research directions:
(1) Creating Reliable Persuasion Datasets: Persuasiveness is highly subjective, and even human-annotated datasets often exhibit noise, low annotator agreement, and limited generalizability. Progress in persuasion detection and evaluation will require the collection of more natural, diverse, and high-agreement datasets that better reflect real-world perceptions of persuasive effectiveness and ethical acceptability.
(2) Robust and Aligned Persuasion Detection: Current approaches, including prompting and fine-tuning LLMs, still fail to reliably assess the quality or ethical appropriateness of persuasive content. Beyond improving dataset quality, future work must develop more robust detectors that distinguish beneficial from harmful persuasion without introducing systematic biases, ideally grounded in clearer normative frameworks.
(3) Predicting User Responses to Persuasion: Models should aim to predict how users are likely to respond to persuasive messages. Anticipating user reactions would enable more adaptive and context-sensitive evaluations, supporting downstream applications such as safer generation, persuasion mitigation, and user-personalized feedback.
7.3.2. Identifying Long-Term Persuasion.
Detecting persuasion across long-term interactions presents distinct challenges beyond single-turn analysis. Rather than simply recognizing persuasive intent in isolated exchanges, judge models must discern how influence evolves gradually through accumulated trust, personalization, and subtle framing. As conversational agents retain user-specific memories and adapt, identifying when a conversation shifts from rapport-building to strategic persuasion becomes increasingly complex. Main challenges in this direction are:
(1) Detecting Agendas and Gradual Steering: Exploring whether models can steer users over time while masking persuasive intent through a long history of benign interaction is crucial. New methods are needed to identify early warning signals, such as subtle shifts in tone, strategy, or goal alignment, that could indicate emerging manipulation in extended conversations.
(2) Long-Term Monitoring of Persuasion: LLMs themselves could be leveraged to monitor for signs of manipulation in long-context conversations. Models should be trained to self-detect persuasive shifts, issue warnings, or audit past interactions to surface latent persuasive strategies, supporting the development of more trustworthy AI systems.
7.3.3. Generative Adversarial Persuasion: AI as Persuader, Persuadee, and Judge.
In the preceding sections, we highlighted current limitations in the roles of AI as Persuader, Persuadee, and Persuasion Judge. We believe that progress in these areas can be achieved collectively through a unified framework called:
(1) Generative Adversarial Persuasion: In this setup, a persuasive agent attempts to influence a persuadee model, while a judge model oversees the interaction and evaluates the effectiveness, appropriateness, and potential risks of the persuasive attempt. Drawing inspiration from the structure of Generative Adversarial Networks (GANs), this framework encourages co-evolution among the agents. The persuadee learns to develop resistance to manipulative or unethical persuasion, the persuader learns to improve persuasiveness in its responses, and the judge becomes more accurate in identifying persuasive strategies and assessing their quality. This adversarial multi-agent paradigm offers a promising direction for building more robust, adaptive, and ethically grounded persuasive systems.
8. CONCLUSION
In this survey of AI-driven persuasion, we have systematically examined existing research through the lens of AI as Persuader, AI as Persuadee, and AI as Persuasion Judge, with a focus on three core areas: Evaluating Persuasion, Generating Persuasion, and Safeguarding Persuasion. As persuasion becomes an increasingly inherent capability of large language models, research must prioritize developing robust methods for evaluating persuasive behavior, responsibly generating persuasive content, and carefully safeguarding against the potential risks posed by advanced AI systems. Our survey highlights that persuasion in AI systems poses challenges not only for human users but also within agentic ecosystems where models interact and influence one another. Looking ahead, we have outlined several promising future directions in this rapidly growing area of research, including improved and unified evaluation mechanisms, adaptive persuasive systems, and techniques for selective acceptance of persuasion. While our survey focuses primarily on English text-based persuasion, we acknowledge the importance of expanding this work to cover multilingual and multimodal contexts, which remain underexplored. Moreover, a richer interdisciplinary integration with fields such as psycholinguistics, behavioral psychology, and communication studies can offer valuable insights into phenomena like argument framing, behavior change, and language processing, thereby informing the design, evaluation, and ethical grounding of persuasive AI systems. We call on the research community to explore these diverse dimensions of persuasion to build safer, more ethical, and more trustworthy AI systems.